A coffee-cup calorimeter contains 50.0 g of water at 60.51°C. A 12.4 g piece of graphite at 24.21°C is  

     placed in the calorimeter. The final temperature of the water and the carbon is 59.02°C. Calculate

     the specific heat of carbon. The specific heat of water is 4.18 J/g⋅°C.


1
Expert's answer
2022-04-21T13:38:04-0400

 

In general, Q = cmΔT

For water: Q1 = 4.18 * 50.0 * (60.51 – 59.02) = 311.41 (J)

For graphite: Q2= c * 12.4 * (59.02 – 24.21)

Q1 = Q2,

c * 12.4 * (59.02 – 24.21) = 311.41

c = 0.72 (J/g⋅°C)